OpenStudy (anonymous):

Two flasks are connected with a stopwingspan. The first flask has a volume of 5 liters and contains nitrogen gas at a pressure of 0.75 atm. The second flask has a volume of 8 L and contains oxygen gas at a pressure of 1.25 atm. When the stopwingspan between the flasks is opened and the gases are free to mix, what will the pressure be in the resulting mixture?

OpenStudy (anonymous):

pV = nRT p = (0.75 * 5 + 1.25 * 8) / (5 +8) = 1.06 atm

OpenStudy (anonymous):

I think that's correct @AccessDenied

OpenStudy (accessdenied):

I think it is correct, although I was going to use Boyle's Law and Dalton's Law instead: P1 V1 = P2 V2 (a) 0.75 atm * 5L = P2 * 13 L P2 = 0.75*5/13 atm (b) 1.25 atm * 8L = P2 * 13L P2 = 1.25*8/13 atm Dalton's Law: P_total = P_O2 + P_N2 = 0.75*5/13 + 1.25*8/13 = (.075*5 + 1.25*8)/13 atm maybe i'm just not used to it but it did not appear obvious from that work to me.

OpenStudy (anonymous):

Great! xD

OpenStudy (anonymous):

Thank you